Write the algorithm, flow chart and the basic programs to complete the sum of the first ten integers.

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Write the algorithm, flow chart and the basic programs to complete the sum of the first ten integer's [sic]
===============
There are 10 of them.
Add 5 pairs:
1 + 10 = 11
2 + 9 = 11
etc.
--> 5*11 = 55
=====================
PS It's integers, not integer's

Answer by Edwin McCravy(20055)


Algorithm:

1. Start with N=0, 
2. Start with SUM=0, 
3. Add 1 to N
4. Add N to SUM
5. If N=10 then print SUM, if not, go to 3
6. End


The following should give you an idea of how
your flow chart should look:

    N=0
     ↓
   SUM=0
     ↓
Add 1 to N ← ←
     ↓        ↑
Add N to SUM  ↑
     ↓        ↑
  N=10? → NO →      <--This is a decision, so
     ↓                 enclose in diamond shape,
    YES                others in rectangles 
     ↓
 Print SUM
     ↓
    End

Here's the program in LibertyBasic.  You may have to 
translate it into whatever language you are using:

  N=0
  SUM=0
1 N=N+1
  SUM=SUM+N
  if N=10 then print SUM else 1
  end

Edwin